OpenStudy (anonymous):

3(4y-7)>8(2y+3) THIS IS UNDERLINED >

OpenStudy (amistre64):

>/ or >= works too

OpenStudy (amistre64):

distribute thru your ()s -12y from the sides and -24 as well

OpenStudy (amistre64):

then divide it al by 4

OpenStudy (anonymous):

-32?

OpenStudy (amistre64):

-45/4 >= y

OpenStudy (anonymous):

improper fraction?

OpenStudy (anonymous):

\[3(4y-7)\geq8(2y+3)\] \[12y-21\geq 16y+24\] \[-21-24\geq 16y-12y\] \[-45\geq 4y\] \[-\frac{45}{4}\geq y\]

OpenStudy (amistre64):

unless you typoed the question; thats what it amounts to

OpenStudy (anonymous):

whew too much cote de rhone

OpenStudy (anonymous):

symboly is the wrong way

OpenStudy (amistre64):

lol..... the symbol doesnt hange the math of it all tho ;)